# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ B.2 Proofs of Section 4
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Define $\mathcal{Z}\coloneqq\operatorname{ran}{g(U+V)}$ and $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Then $\dim{\mathcal{Z}}<\infty$ and $\operatorname{ran}{g(U+tV)}\subset\mathcal{Z}$ for all $t\in\mathbb{R}$, hence $g(U+tV)\in\mathcal{W}$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g(U+tV)={\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)$ for all $t\in\mathbb{R}$. By Lemma 4.9 and Proposition 4.14, ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable. By Lemma 4.9, $g$ is Fréchet differentiable. In particular, $g$ is Gateaux differentiable in $U$ in direction $V$. Hence, by the chain rule, $J_{f}$ is Gateaux differentiable in $U$ in direction $V$. To compute the derivative, we recall that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. By the expression of the derivative in Proposition 4.14, the Gateaux derivative of $J_{f}$ in $U$ in direction $V$ is
@@
$=\langle{\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(g(U)),g^{\prime}(U)V\rangle_{L_{2}(\mathcal{H})}$
$=\langle\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},g^{\prime}(U)V\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{j,k}\langle\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},e_{j}\otimes e_{k}\rangle_{L_{2}(\mathcal{H})}\langle g^{\prime}(U)V,e_{j}\otimes e_{k}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle g^{\prime}(U)V,e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2},e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}$
@@
Let $f\in\mathscr{F}$ and let $(U_{n})_{n}$ be a sequence in $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ such that $\lVert U_{n}\rVert\rightarrow\infty$. Then, by Lemma A.3, $\lVert U_{n}U_{n}^{*}\rVert=\lVert U_{n}\rVert^{2}\rightarrow\infty$. Since $f(x)\rightarrow\infty$ for $\lVert x\rVert\rightarrow\infty$ and $f$ is bounded from below, it is enough to show that there is an eigenvalue $\alpha_{n}$ of $g(U_{n})$ with $\lvert\alpha_{n}\rvert\rightarrow\infty$. Since $g(U_{n})$ is self-adjoint and compact, $\lVert g(U_{n})\rVert=\max_{i}\lvert\Lambda_{i}(g(U_{n}))\rvert$ by Lemma A.4, and we must therefore show that $\lVert g(U_{n})\rVert\rightarrow\infty$. For this, it is enough to find a bounded sequence $h_{n}\in\mathcal{H}$, such that $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$.
For any $h\in\mathcal{H}$, we have by the triangle inequality, (18) and Lemma 3.6
$\lVert g(U)h\rVert$ $=\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h-(I-(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})h\rVert$
$\geq\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rVert-\lVert(I-(\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})^{*}\mathcal{C}_{\textup{pr}}^{-1/2}\mathcal{C}_{\textup{pos}}^{1/2})h\rVert$
@@
Let us write $m\coloneqq\dim{\mathcal{V}}$. For each $n$, let us diagonalise $U_{n}U_{n}^{*}=\sum_{j=1}^{m}\beta_{n,j}\psi_{n,j}\otimes\psi_{n,j}$, where $(\psi_{n,j})_{j=1}^{m}$ forms an ONB of $\mathcal{V}$ and $\beta_{n,j}\geq 0$. Define $j_{n}\coloneqq\operatorname*{arg\,max}_{j\leq m}\beta_{n,j}$, hence $\beta_{n,j_{n}}$ is the largest eigenvalue of $U_{n}U_{n}^{*}$. As $U_{n}U_{n}^{*}$ is self-adjoint, $\beta_{n,j_{n}}=\lVert U_{n}U_{n}^{*}\rVert$ by Lemma A.4, showing $\beta_{n,j_{n}}\rightarrow\infty$. Let $\varepsilon>0$. By density of $\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$, for each $j\leq m$ we can choose $k_{j}\in\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$ which satisfies $\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert\leq\varepsilon$. Let us decompose $\psi_{n,j_{n}}=\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\psi_{1,j}$ and define $h_{n}\coloneqq\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle k_{j}$. Note that $\lVert h_{n}\rVert\leq C$ with $C\coloneqq m\max_{j}\lVert k_{j}\rVert$ by the Cauchy–Schwarz inequality. By further application of the Cauchy–Schwarz inequality,
@@
Thus, by the definition of $(k_{j})_{j}$, the triangle inequality and the bound $(a+b)^{2}\leq 2a^{2}+2b^{2}$,
$\lVert\mathcal{C}_{\textup{pos}}^{1/2}k_{n}\rVert^{2}\geq\frac{1}{2}\lVert\psi_{n,j_{n}}\rVert^{2}-\frac{1}{2}\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}\geq\frac{1}{2}-m(m-\frac{1}{2})\varepsilon^{2}.$
@@
$\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle=\frac{1}{2}\left(\lVert\psi_{n,j_{n}}\rVert^{2}+\lVert\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}-\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}\right)>c.$
@@
$=\sum_{j}\beta_{n,j}\lvert\langle\psi_{n,j},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq C^{-1}\beta_{n,j_{n}}\lvert\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq c^{2}C^{-1}\beta_{n,j_{n}}\rightarrow\infty.$
Combining this with (44), we have thus found a bounded sequence $(h_{n})_{n}$ with $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$. Finally, by Proposition 4.15, $J_{f}$ is differentiable. A coercive, differentiable function has a global minimum, that is attained only among its stationary points. ∎
See 4.18
@@
With $(\gamma_{i})_{i}$ and $(e_{i})_{i}$ as in Lemma 4.7, the Gateaux derivative $J_{F}^{\prime}(g(U))\in L_{2}^{*}\simeq L_{2}$ at $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is given by, in direction $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$,
$J_{f}^{\prime}(U)(V)=2\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle.$
Setting the derivative to zero, we obtain for all $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$,

### v3

@@ B.2 Proofs of Section 4
Any statement regarding differentiability in this proof should be understood as Fréchet differentiability. For $Y\in\mathcal{W}$, we have $\ker{Y}^{\perp}=\ker{Y^{*}}^{\perp}=\overline{\operatorname{ran}{Y}}=\operatorname{ran}{Y}$, by Lemma A.6, $Y=Y^{*}$ and the finite dimensionality of $\mathcal{Z}$. Let $m\coloneqq\dim{\mathcal{Z}}$ and extend $(e_{i})_{i}$ to an ONB of $\mathcal{H}$. Note that the $m^{2}$ operators $e_{i}\otimes e_{j}$, $i,j\leq m$, span the space $\mathcal{W}$. Therefore, $\dim{\mathcal{W}}=m^{2}<\infty$. Because finite-dimensional spaces are closed, we can define $P_{\mathcal{Z}}:\mathcal{H}\rightarrow\mathcal{Z}$, the orthogonal projector onto $\mathcal{Z}$. Furthermore, we let $P_{m}:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}^{m}$ be the orthogonal projector onto the first $m$ coordinates of an $\ell^{2}$ sequence. Thus, $P_{m}^{*}$ is the natural embedding of $\mathbb{R}^{m}$ into $\ell^{2}(\mathbb{R})$. Since Proposition 4.14 is a statement on $L_{2}$ spaces of finite dimension, we identify $\mathcal{W}$ with $L_{2}(\mathcal{Z})_{\mathbb{R}}$ via $\Psi:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow L_{2}(\mathcal{Z})_{\mathbb{R}}$, where $\Psi(Y)\coloneqq P_{\mathcal{Z}}{\left.\kern-1.2ptY\vphantom{\big{|}}\right|_{\mathcal{Z}}}$.
We first prove the result for the case in which $\Lambda$ orders the eigenvalues in nonincreasing absolute value. With $\Lambda^{m}$ denoting the eigenvalue map on $L_{2}(\mathcal{Z})_{\mathbb{R}}$ as defined in Section 1.5, we then have $\Lambda(Y)_{i}=\Lambda^{m}(\Psi(Y))_{i}$ for all $i\leq m$ and $Y\in\mathcal{W}$. This implies that
$P_{m}\Lambda(Y)$ $=\Lambda^{m}(\Psi(Y)),\quad Y\in\mathcal{W},$
$\Lambda(Y)$ $=P_{m}^{*}\Lambda^{m}(\Psi(Y)),\quad Y\in\mathcal{W},$ (27)
since any $Y\in\mathcal{W}$ has at most $m$ nonzero eigenvalues. Let $\mathcal{G}\coloneqq FP_{m}^{*}$. If $\pi$ is a permutation on $\{1,\ldots,m\}$ and $x\in\mathbb{R}^{m}$, then $P_{m}^{*}x=(x_{1},\ldots,x_{m},0,\ldots)$ and $P_{m}^{*}(x_{\pi(i)})_{i}=(x_{\pi(1)},\ldots,x_{\pi(m)},0,\ldots)$, and by symmetry of $F$,
$\mathcal{G}((x_{\pi(i)})_{i=1}^{m})=F(P_{m}^{*}(x_{\pi(i)})_{i})=F(P_{m}^{*}x)=\mathcal{G}(x)$
showing that $\mathcal{G}$ is a symmetric function on the symmetric domain $\mathbb{R}^{r}$. Furthermore, by definition of $\mathcal{G}$ and (27)
${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}(Y)=(F\circ\Lambda)(Y)=(\mathcal{G}\circ\Lambda^{m})(\Psi(Y)),\quad Y\in\mathcal{W}.$ (28)
By hypothesis, $F$ is differentiable at $\Lambda(X)$, with $X\in\mathcal{W}$. The idea of the proof is to first use Proposition 4.14 to conclude differentiability of $\mathcal{G}\circ\Lambda^{m}$ at $\Psi(X)$ and then to use (28) and the chain rule to obtain differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ at $X$ in the $L_{2}(\mathcal{H})$ norm topology.
In order to apply Proposition 4.14, we need to show that $\mathcal{G}$ is differentiable in $\Lambda^{m}(\Psi(X))$. By the hypothesis on $F$, $F$ is differentiable at $\Lambda(X)$ for $X\in\mathcal{W}$. Furthermore, $P_{m}^{*}$ is linear, hence differentiable, and $(P_{m}^{*})^{\prime}(x)(y)=P_{m}^{*}y$ for $x,y\in\mathbb{R}^{m}$. Then, by (27) and the chain rule, the composition $F\circ P_{m}^{*}$ is differentiable at $\Lambda^{m}(\Psi(X))$, and it holds for any $y\in\mathbb{R}^{m}$,
$\langle(FP_{m}^{*})^{\prime}(\Lambda^{m}(\Psi(X))),y\rangle$ $=\langle F^{\prime}(P_{m}^{*}\Lambda^{m}(\Psi(X))),(P_{m}^{*})^{\prime}(\Lambda^{m}(\Psi(X)))y\rangle$
$=\langle F^{\prime}(P_{m}^{*}\Lambda^{m}(\Psi(X))),P_{m}^{*}y\rangle$
$=\langle P_{m}F^{\prime}(P_{m}^{*}\Lambda^{m}(\Psi(X))),y\rangle$
$=\langle P_{m}F^{\prime}(\Lambda(X)),y\rangle,$
where we use the chain rule in the first step, the expression for the derivative $(P_{m}^{*})^{\prime}$ in the second step, the definition of the adjoint in the third step, and (27) in the final step. That is, $\mathcal{G}$ is differentiable at $\Lambda^{m}(\Psi(X))$ and
$\mathcal{G}^{\prime}(\Lambda^{m}(\Psi(X)))=P_{m}F^{\prime}(\Lambda(X))\in\mathbb{R}^{m}.$ (29)
We may now apply Proposition 4.14 to conclude that $\mathcal{G}\circ\Lambda^{m}$ is differentiable at $\Psi(X)$. To obtain an expression for the derivative, notice that by the fact that $e_{i}\in\operatorname{ran}{\mathcal{Z}}$ for $i\leq m$ and by the hypothesised diagonalisation of $X$, we have $\Psi(X)=\sum_{i=1}^{m}\Lambda^{m}(\Psi(X))_{i}e_{i}\otimes e_{i}$, where the rank-1 operators $e_{i}\otimes e_{i}$ are now understood to act only on $\mathcal{Z}$. With Proposition 4.14 we thus also obtain the expression for the derivative
$(\mathcal{G}\circ\Lambda^{m})^{\prime}(\Psi(X))=\sum_{i=1}^{m}\mathcal{G}^{\prime}(\Lambda^{m}(\Psi(X)))_{i}e_{i}\otimes e_{i}=\sum_{i=1}^{m}\left(P_{m}F^{\prime}(\Lambda(X))\right)_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z})_{\mathbb{R}},$
where for second equation we use (29). By definition of $P_{m}$, $\left(P_{m}F^{\prime}(\Lambda(X))\right)_{i}=F^{\prime}(\Lambda(X))_{i}$ for $i\leq m$. Hence,
$(\mathcal{G}\circ\Lambda^{m})^{\prime}(\Psi(X))=\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i}.$ (30)
Because $\Psi$ is linear, hence differentiable, the chain rule and (28) show that ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable at $X$. To obtain the expression of the derivative, we use (28), the chain rule, the fact $\Psi^{\prime}(X)(Y)=\Psi(Y)$ for $Y\in\mathcal{W}$ and (30) to find
$\langle{\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(X),Y\rangle_{L_{2}(\mathcal{H})}$ $=\langle(\mathcal{G}\circ\Lambda^{m}\circ\Psi)^{\prime}(X),Y\rangle_{L_{2}(\mathcal{H})}=\langle(\mathcal{G}\circ\Lambda^{m})^{\prime}(\Psi(X)),\Psi^{\prime}(X)(Y)\rangle_{L_{2}(\mathcal{Z})}$
$=\langle(\mathcal{G}\circ\Lambda^{m})^{\prime}(\Psi(X)),\Psi(Y)\rangle_{L_{2}(\mathcal{Z})}=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},\Psi(Y)\Big{\rangle}_{L_{2}(\mathcal{Z})}.$
Since, for $Y\in\mathcal{W}$, it holds that $\operatorname{ran}{Y}\subset\mathcal{Z}$ and $\ker{Y}^{\perp}=\operatorname{ran}{Y}$, we have $\ker{Y}^{\perp}\subset\mathcal{Z}$. Thus, we have $\operatorname{ran}{\Psi(Y)}=Y(\mathcal{Z})=Y(\ker{Y}^{\perp})=\operatorname{ran}{Y}\subset\mathcal{Z}$ as subspaces of $\mathcal{H}$. For $i\leq m$ it holds that $e_{i}\in\mathcal{Z}$, hence $\langle e_{i}\otimes e_{i},\Psi(Y)\rangle_{L_{2}(\mathcal{Z})}=\langle e_{i}\otimes e_{i},Y\rangle_{L_{2}(\mathcal{H})}$, where on the right hand side we interpret $e_{i}\otimes e_{i}$ as acting on all of $\mathcal{H}$. For $i>m$, we have $e_{i}\in\mathcal{Z}^{\perp}$, so that $\operatorname{ran}{Y}\subset\mathcal{Z}$ implies that $\langle e_{i}\otimes e_{i},Y\rangle_{L_{2}(\mathcal{H})}=0$. Thus,
$\langle{\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(X),Y\rangle_{L_{2}(\mathcal{H})}=\Big{\langle}\sum_{i=1}^{m}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})}=\Big{\langle}\sum_{i=1}^{\infty}F^{\prime}(\Lambda(X))_{i}e_{i}\otimes e_{i},Y\Big{\rangle}_{L_{2}(\mathcal{\mathcal{H}})}.$
This concludes the proof for the case that $\Lambda$ orders the eigenvalues in a nonincreasing way.
Finally, let us denote by $\tilde{\Lambda}$ an eigenvalue map on $L_{2}(\mathcal{H})_{\mathbb{R}}$ which can assign any fixed but arbitrary ordering on the eigenvalues. Let $X\in\mathcal{W}$, $X=\sum_{i=1}\tilde{\Lambda}(X)_{i}e_{i}\otimes e_{i}$ be given and assume that $F$ is differentiable at $\tilde{\Lambda}(X)$. Given $X$, there exists a permutation $\pi:\mathbb{N}\rightarrow\mathbb{N}$ such that, for the eigenvalue map $\Lambda$ from the previous part of the proof, $\tilde{\Lambda}(X)_{i}=\Lambda(X)_{\pi(i)}$ for all $i\in\mathbb{N}$. Let $P_{\pi}:\ell^{2}(\mathbb{R})\rightarrow\ell^{2}(\mathbb{R})$ denote the permutation operator $(P_{\pi}x)_{i}=x_{\pi(i)}$, $i\in\mathbb{N}$, so $P_{\pi}^{*}=P_{\pi}^{-1}$. Then $\tilde{\Lambda}(X)=(P_{\pi}\Lambda)(X)$ and $X=\sum_{i}\Lambda_{\pi(i)}(X)e_{i}\otimes e_{i}=\sum_{i}\Lambda(X)_{i}e_{\pi^{-1}(i)}\otimes e_{\pi^{-1}(i)}.$ By the previous part of the proof, ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is differentiable at $X$. Because $F\circ\tilde{\Lambda}=F\circ\Lambda$ by symmetry of $F$, differentiability of ${\left.\kern-1.2pt(F\circ\tilde{\Lambda})\vphantom{\big{|}}\right|_{\mathcal{W}}}$ at $X$ follows, and
$(F\circ\tilde{\Lambda})^{\prime}(X)=(F\circ\Lambda)^{\prime}(X)=\sum_{i}F_{i}^{\prime}(\Lambda(X))e_{\pi^{-1}(i)}\otimes e_{\pi^{-1}(i)}.$ (31)
Since $F$ is symmetric, $F\circ P_{\pi}=F$. Hence, for $x\in\ell^{2}(\mathbb{R})$,
$F^{\prime}(x)=(F\circ P_{\pi})^{\prime}(x)=P_{\pi}^{*}F^{\prime}(P_{\pi}x)=P_{\pi}^{-1}F^{\prime}(P_{\pi}x).$
Thus, $F_{i}^{\prime}(\Lambda(X))=F_{\pi^{-1}(i)}^{\prime}(P_{\pi}\Lambda(X))=F_{\pi^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))$. From (31) it now follows that
$(F\circ\tilde{\Lambda})^{\prime}(X)=\sum_{i}F_{\pi^{-1}(i)}^{\prime}(\tilde{\Lambda}(X))e_{\pi^{-1}(i)}\otimes e_{\pi^{-1}(i)}=\sum_{i}F_{i}^{\prime}(\tilde{\Lambda}(X))e_{i}\otimes e_{i}.$
See 4.18
Proof of Proposition 4.18.
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Define $\mathcal{Z}\coloneqq\operatorname{ran}{g(U+V)}$ and $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Then $\dim{\mathcal{Z}}<\infty$ by Lemma 4.10 and $\operatorname{ran}{g(U+tV)}\subset\mathcal{Z}$ for all $t\in\mathbb{R}$ by definition (19) of $g$, hence $g(U+tV)\in\mathcal{W}$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g(U+tV)={\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)$ for all $t\in\mathbb{R}$. By Lemma 4.11 and Proposition 4.16, ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable. By Lemma 4.11, $g$ is Fréchet differentiable. In particular, $g$ is Gateaux differentiable in $U$ in direction $V$. Hence, by the chain rule, $J_{f}$ is Gateaux differentiable at $U$ in the direction $V$. To compute the derivative, we recall that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. The Gateaux derivative of $J_{f}$ at $U$ in the direction $V$ is
@@
$=\Big{\langle}{\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}^{\prime}(g(U)),g^{\prime}(U)(V)\Big{\rangle}_{L_{2}(\mathcal{H})}$
$=\Big{\langle}\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},g^{\prime}(U)(V)\Big{\rangle}_{L_{2}(\mathcal{H})}$
$=\sum_{j,k}\Big{\langle}\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)e_{i}\otimes e_{i},e_{j}\otimes e_{k}\Big{\rangle}_{L_{2}(\mathcal{H})}\langle g^{\prime}(U)(V),e_{j}\otimes e_{k}\rangle_{L_{2}(\mathcal{H})}$
$=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle g^{\prime}(U)(V),e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}.$
The second equation follows by the chain rule. The third equation follows from the expression for the derivative in Proposition 4.16 applied with $X\leftarrow g(U)$ and the expression for $F_{f}^{\prime}$ in Lemma 4.11. The fourth and fifth equations use the property that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. Using the formula for $g^{\prime}(U)(V)$ from Lemma 4.11,
$J_{f}^{\prime}(U)(V)$ $=\sum_{i}f^{\prime}\left(\Lambda_{i}(g(U))\right)\langle\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2},e_{i}\otimes e_{i}\rangle_{L_{2}(\mathcal{H})}$
@@
Let $f\in\mathscr{F}$ and let $(U_{n})_{n}$ be a sequence in $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ such that $\lVert U_{n}\rVert\rightarrow\infty$. Then, by Lemma A.2, $\lVert U_{n}U_{n}^{*}\rVert=\lVert U_{n}\rVert^{2}\rightarrow\infty$. Since $f(x)\rightarrow\infty$ for $\lVert x\rVert\rightarrow\infty$ and $f$ is bounded from below, it is enough to show that there is an eigenvalue $\alpha_{n}$ of $g(U_{n})$ with $\lvert\alpha_{n}\rvert\rightarrow\infty$. Since $g(U_{n})$ is self-adjoint and compact by definition (19), $\lVert g(U_{n})\rVert=\max_{i}\lvert\Lambda_{i}(g(U_{n}))\rvert$ by Lemma A.3, and we must therefore show that $\lVert g(U_{n})\rVert\rightarrow\infty$. For this, it is enough to find a bounded sequence $h_{n}\in\mathcal{H}$, such that $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$.
For any $h\in\mathcal{H}$, we have by the triangle inequality, (19) and Proposition 3.7
$\lVert g(U)h\rVert$ $=\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h-\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}h\rVert$
$\geq\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}h\rVert-\lVert\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}h\rVert$
@@
Let us write $m\coloneqq\dim{\mathcal{V}}$. For each $n$, let us diagonalise $U_{n}U_{n}^{*}=\sum_{j=1}^{m}\beta_{n,j}\psi_{n,j}\otimes\psi_{n,j}$, where $(\psi_{n,j})_{j=1}^{m}$ forms an ONB of $\mathcal{V}$ and $\beta_{n,j}\geq 0$. Define $j_{n}\coloneqq\operatorname*{arg\,max}_{j\leq m}\beta_{n,j}$, hence $\beta_{n,j_{n}}$ is the largest eigenvalue of $U_{n}U_{n}^{*}$. As $U_{n}U_{n}^{*}$ is self-adjoint, $\beta_{n,j_{n}}=\lVert U_{n}U_{n}^{*}\rVert$ by Lemma A.3, showing $\beta_{n,j_{n}}\rightarrow\infty$. Let $\varepsilon>0$. By density of $\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$, for each $j\leq m$ we can choose $k_{j}\in\mathcal{H}$ which satisfies $\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert\leq\varepsilon$. Let us decompose $\psi_{n,j_{n}}=\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\psi_{1,j}$ and define $h_{n}\coloneqq\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle k_{j}$. Note that $\lVert h_{n}\rVert\leq C$ with $C\coloneqq m\max_{j}\lVert k_{j}\rVert$ by the Cauchy–Schwarz inequality. By further application of the Cauchy–Schwarz inequality,
@@
$\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle=\frac{1}{2}\left(\lVert\psi_{n,j_{n}}\rVert^{2}+\lVert\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}-\lVert\psi_{n,j_{n}}-\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rVert^{2}\right)\geq\frac{1}{2}(1+0-m(2m-1)\varepsilon^{2})>c.$
@@
$=C^{-1}\sum_{j}\beta_{n,j}\lvert\langle\psi_{n,j},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq C^{-1}\beta_{n,j_{n}}\lvert\langle\psi_{n,j_{n}},\mathcal{C}_{\textup{pos}}^{1/2}h_{n}\rangle\rvert^{2}\geq c^{2}C^{-1}\beta_{n,j_{n}}\rightarrow\infty.$
Combining this with (32), we have thus found a bounded sequence $(h_{n})_{n}$ with $\lVert g(U_{n})h_{n}\rVert\rightarrow\infty$. Finally, by Proposition 4.18, $J_{f}$ is differentiable. The conclusion follows because a coercive, differentiable function on a finite-dimensional space $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ has a global minimum, that is attained only among its stationary points. ∎
See 4.21
@@
Let $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Let $(\gamma_{i})_{i}$ and $(e_{i})_{i}$ be as in Lemma 4.10, so that $e_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$ and $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$. By Proposition 4.18, the Gateaux derivative $J_{f}^{\prime}(g(U))\in L_{2}^{*}\simeq L_{2}$ at $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ is given by
$J_{f}^{\prime}(U)(V)=2\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle,\quad V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H}).$
Thus, $U$ is a stationary point of $J_{f}$ if and only if for all $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$,